OpenStudy (anonymous):

Please find the exact value of the radical expression in simplest form. √4y^3 - √y^3

OpenStudy (anonymous):

\[\large \sqrt{4y^3} - \sqrt{y^3} \implies 2y \sqrt{y} - y \sqrt{y} \implies \color{green}{y \sqrt{y}}\]